OpenStudy (ledah):

Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5).

OpenStudy (anonymous):

first solve for y and you will get something in y=mx+b form the m will be the same in both your equations.

OpenStudy (ledah):

y = -4x + 20 like this? this is what i got

OpenStudy (anonymous):

y = -2*x+5/2 y= m x+ b so your m will be -2 so now plug in your values for the point you want (-3,5) 5=-2*-3+b sfb -1 so your equation is y=-2x+(-1)

OpenStudy (anonymous):

nope it is not y=-4....

OpenStudy (ledah):

oh. oops fail lol

OpenStudy (anonymous):

4x + 2y = 5 2y=5-4x subtract 4x y=-2x+5/2 rearange and subtract 2 from both sides